What multiplies to be -6 but adds to be 7?

Answer by stanbon(75887)

What multiplies to be -6 but adds to be 7?

---------------------
Equations:
x*y = -6
x + y = 7
------------------
Substitute for "x" and solve for "y"::
(7-y)y = -6
----
-y^2 + 7y = -6
---
y^2 - 7y + -6 = 0
-----
y = [7 +- sqrt(49-4*-6)]/2
-----
y = [7+- sqrt(73)]/2
y = 3.5 + (1/2)sqrt(73) or y = (3.5)-(1/2)sqrt(73)
----------------
Solve for "x" = 7-y ::
If y = 3.5 + (1/2)sqrt(173), x = 10.5-(1/2)sqrt(173)
If y = 3.5 - (1/2)sqrt(173), x = 10.5+(1/2)sqrt(173)
=============
Cheers,
Stan H.
-------------